OpenStudy (anonymous):

Find the solution to the equation log16 32 = x + 5.

OpenStudy (anonymous):

\[\log_{16} (32)=x+5\]

OpenStudy (nurali):

if 16^(x+5) = 32

OpenStudy (anonymous):

is that your question?

OpenStudy (nurali):

take logs of both sides: log 16^(x+5) = log 32 by laws of logarithms log a^x = x log a: (x+5) log 16 = log 32 x+5 = log 32/log 16 = 1.25 x = 1.25 - 5 = -3.75

OpenStudy (anonymous):

Yes ! @ASAAD123